Find an equation of the tangent plane to the following parametric surface, r(u, v)  =  (u2 + 9) i + (v3 + 6u) j + (u + 2v) k , at the point (10, 5, −1). Write the equation in the form ax + by + cz + d  =  0, where a, b, c, and d have no common factors. Then enter the values of a, b, c, and d (in that order) into the answer box below, separated with commas
Answer #1

wote: please give good Thumb up which helps helps me to f of Answer my more do answersHere are the given ruu) = (479)i + (W²+64)J +14+20) ī Lets first compute ru xro. two individual vectors rucu,o) qui tout 201qi -uJ + GF : required tangent plane equation is 968-10) - 4 (7-5)+617 +1)=0 =) qx-90-49 +20 +62 +6=0 =) 94-444 62-64=0 ia=9,